# Evolution of Photonic Quantum Machine Learning under Noise

### 10 Mar 2026 v1

@@ Abstract
Photonic Quantum Machine Learning (PQML) is an emerging method to implement scalable, energy-efficient quantum information processing by combining photonic quantum computing technologies with machine learning techniques. The features of photonic technologies offer several benefits: room-temperature operation; fast (low delay) processing of signals; and the possibility of representing computations in high-dimensional (Hilbert) spaces. This makes photonic technologies a good candidate for the near term development of quantum devices. However, noise is still a major limiting factor for the performance, reliability, and scalability of PQML implementations. This review provides a detailed and systematic analysis of the sources of noise that will affect PQML implementations. We will present an overview of the principal photonic quantum computer designs and summarize the many different types of quantum machine learning algorithms that have been successfully implemented using photonic quantum computer architectures as Variational Quantum Circuits, Quantum Neural Networks, and Quantum Support Vector Machines. We identify and categorize the primary sources of noise within photonic quantum systems, and how these sources of noise behave algorithm-specifically with respect to degrading the accuracy of learning, unstable training, and slower convergence than expected. Additionally, we review traditional and advanced techniques for characterizing noise, and provide an extensive survey of strategies for mitigating the effects of noise on learning performance. Finally, we discuss recent advances that demonstrate PQML’s capability to operate in real-world settings with realistic noise conditions and future obstacles that will challenge the use of PQML as an effective quantum processing platform.
Preprint. This manuscript is currently under review at Quantum Machine Intelligence (Springer Nature).
@@ 1 Introduction
Quantum computing utilizes the principles of quantum mechanics such as superposition and entanglement to solve problems that the classical computer cannot. First ever ”universal quantum computer” idea was published in 1985 by David Deutsch [17]. In 1994, Peter Shor presents an algrithm that can efficiently find the factors of large numbers, which is latter known as Shor’s Algorithm [43]. When you perform a computation on quantum bits (qubits), which can be in multiple states at once due to superposition, this allows for parallel processing of that computation. Qubits are also intertwined or entangled, which results in them being correlated in a strong way, which creates very strong interactions and increases the computer’s power[20]. Quantum computers can revolutionize machine learning by introducing algorithms that run much faster, and can process the data more efficiently.
Figure 1: Timeline of major advancements in quantum machine learning [6].
@@
Quantum Machine Learning (QML) is a multidiscipline that brings together quantum computer science and traditional machine learning. There is no exact date marking the beginning of QML. But Subhash C. Kak published a paper with idea about a biologically inspired neural network works on quantum machine learning in 1995 [23]. “Figure. 1” shows the timeline of the major advancements in quantum machine learning. QML seeks to harness the special nature of quantum mechanics to make machine learning more efficient and effective. Quantum computers may have the potential to outperform classical machine learning models for certain tasks by exploiting quantum parallelism and other quantum phenomena. QML currently possesses limitations including noise in quantum systems, limited qubit counts, and the requirement for specialized hardware. However, there is ongoing research into these challenges, and QML has potential to make great advancements to the machine learning field [18].
@@
Among the various quantum computing platforms, photonic quantum computing has attracted particular interest due to the possibility for operation at room temperature [52], and that photonic systems allow low-latency real-world quantum information processing. In photonic quantum computing, qubits are denoted by photons which is the basic unit of light which can be processed by optical components like beam splitters, phase shifters, mirrors, and others to perform quantum operations. Photonic Quantum Machine Learning (PQML) combines photonic systems with quantum learning algorithms, and provides an attractive methodology for potential scalable energy-efficient, low resource, high-fidelity quantum computation [3]. “Figure. 2” illustrates the main components of a photonic quantum computing system.
Figure 2: Illustration of the photonic quantum computing components [1].
@@
In this review, we explore the impact of noise in photonic quantum machine learning systems. The upcoming sections of the paper are organised as follows. Section 2 discusses the fundamentals of photonic quantum machine learning. Section 3 identifies and categorizes the various noise sources in photonic quantum systems. Section 4 analyses the impact of noise on quantum machine learning algorithms. Section 5 reviews noise characterization methods. Section 6 explores noise mitigation strategies. Section 7 highlights recent advances and experimental demonstrations. Section 8 provides a discussion, where theoretical frameworks and noise models are considered in the context of recent experimental demonstrations, along with near-term applications and long-term prospects. Finally, Section 9 concludes the review.
@@ 2.1 Photonic Quantum Computing Architectures
There are several architectures for photonic quantum computing, each with distinct noise characteristics and mitigation requirements. But there is only three main architectures that are widely used in photonic quantum computing. These are discrete-variable (DV) quantum computing, continuous-variable (CV) quantum computing, and hybrid quantum computing.
Figure 3: Single photon on a beam splitter [39].
@@
Continuous Variable (CV) Quantum Computing encodes quantum information in continuous degrees of freedom of the infinite-dimensional Hilbert space [29]. Rather than using discrete qubits, CV quantum computers utilize properties of light modes, such as amplitude and phase, to encode and process information. By leveraging Gaussian operations alongside non-Gaussian resources, CV quantum computing allows for universal quantum computation exploiting optical systems[10]. Noise sources in CV systems include:
@@
Figure 4: Building blocks of a Hybrid Quantum System: Interfacing continuous-variable (EM field, mechanical oscillator) and discrete-variable (spin ensemble) components [7].
Hybrid Quantum Computing amalgamates the two approaches of DV and CV to capitalize on the benefits of both. In hybrid systems we can use discrete qubits encoded in single photons, and CV variables for some parts of the measurement or processing. Noise sources in hybrid systems will include DV and CV noise sources, as well as additional complexities produced by the interaction of the discrete and continuous variables [52]. “Figure. 4” illustrates the building blocks of a hybrid quantum system.
@@ 2.2 Quantum Machine Learning Algorithms on Photonic Platforms
Various quantum machine learning algorithms have been adapted to be used on photonic platforms. Some of the prominent algorithms such as Variational Quantum Circuits (VQC), Quantum Neural Networks (QNN), and Quantum Support Vector Machines (QSVM) have all been successfully implemented in photonic systems [34].So, let’s discuss these algorithms briefly.
Variational Quantum Circuits (VQC) are a powerful framework for quantum machine learning that combines classical optimization techniques and the training of quantum circuits. “Figure. 5” illustrates a basic structure of a VQC. A photonic implementation of VQC can use the huge range of optical components to create tunable quantum circuits, which further allows VQC to successfully train and optimize quantum models [46].
Quantum Neural Networks (QNN) comprise another category of algorithm that can successfully be implemented on photonic platforms. QNNs use quantum circuits for computations similar to classical neural networks, thus operating in the quantum information domain [26]. A key advantage of photo QNNs is the inherent parallelism of photonic states, as well as their high-dimensionality, which benefits learning.
Quantum Support Vector Machines (QSVM) are the quantum counterpart of classical support vector machines to be incorporated to handle tasks of classification with the help of quantum states and operations. In the photonic setting, QSVMs can take advantage of the high-dimensional Hilbert space of photons when encoding data and perform kernel evaluations based on quantum interference effects [4].
Figure 5: A basic diagram of variational quantum circuits [42].
@@ 3 Noise Sources in Photonic Quantum Systems
Photonic quantum systems are vulnerable to different sources of noise that degrade any quantum machine learning algorithm that we are conducting [13]. Recognizing and categorizing these sources of noise is paramount to developing suitable noise mitigation strategies. In this section, we discuss the major sources of noise, broken down into two subsections such as fundamental noise mechanisms and system-specific noise characteristics.
@@ 3.1 Fundamental Noise Mechanisms
While photons are excellent carriers of quantum information, they are still subject to various kinds of noise during generation, transmission, and detection, which can degrade the quality of quantum states and diminish the reliability of practitioners’ quantum operations [21]. Below, we identify the main types of noise in photonic quantum systems, and how these types of noise affect the overall performance of a system.
@@
Photon loss can have a detrimental effect on quantum machine learning algorithms. As an example, in variational quantum circuits, photon loss can result in an erroneous adjustment of parameters while training, and exhibit poor model performance [13]. With respect to quantum neural networks, photon loss can sever the flow of information in the network, degrading learning capabilities [2].
@@ 3.2 System-Specific Noise Characteristics
The other type of noise sources is related to the specific design of a particular photonic quantum device, and the way that the device itself is manufactured like Architectures and Components will also affect how the specific types of noise behave. In this subsection, many examples of photonic quantum device-specific noise have been provided.
  • Mode Mismatch: Mismatched optical modes often cause reduced visibility of interference which lower the values of interference[48]. Also increased error rates on quantum algorithms that rely on interference like boson sampling, quantum walks etc.
@@
As quantum machine learning models rely heavily on the presence of high visibility interference, mode mismatch is one of the most detrimental sources of noise in photonic implementations. The next section will provide a thorough discussion of how mode mismatch and other types of noise represent one of the major sources of impact on photonic implementations from a practical standpoint.
@@ 4.1 Algorithm-Specific Noise Sensitivity
Noise can have a significant impact on the operation of VQC’s and QNN’s. Due to their reliance on the accurate measurement of expectation values and calculation of gradients, variational quantum circuits (VQC’s) are especially vulnerable to the presence of noise [16]. The impact of noise on VQC can be affect in different ways such as:
  • Gradient Vanishing/Exploding: Noise can lead to ”vanishing” or ”exploding” gradients during the optimizational phase making it hard to develop a good quantum model because the gradient is either close to zero or is extreme large. [33]
@@
QNN’s operate using Quantum Mechanics and can be easily impacted by noise during their operation. Therefore, noise may impact the ability of QNN’s to learn and adapt. Similarly, feed forward structures, convolutional and recurrent QNN’s will each experience issues with noise at various levels. So we have to use different strategies to mitigate these noise effects.
@@ 4.2 Noise-Resilient Algorithm Design
Noise -resilient algorithm design focuses on creating quantum machine learning algorithms that can withstand the presence of noise in photonic quantum systems. According to recent research, we can get mainly two types as Hardware aware compilation and algorithmic error supression. “Figure. 6” illustrates the concept of quantum error mitigation (QEM) which aimss to design noise resilient algorithms.
Figure 6: Schematic illustration: (a) noise biases the expected value of an observable in quantum computation, while (b) QEM corrects this bias to recover the ideal expectation value. [38].
Hardware-aware compiling tools for modern quantum circuits take advantage of the specific types of noise present in the underlying hardware to create quantum circuits that are as close to optimal as possible. Specifically, the optimized circuit designs reduce the number of overall errors based on the type of noise present in the photonic platform used. The optimisations include reducing circuit depths and limiting the number of entangling operations within a particular circuit design to reduce the negative effect of photon loss and decoherence.
@@ 5.1 Traditional Noise Characterization Approaches
Figure 7: Comparison of noise spectra reconstructed using FTNS and CPMG-based methods. [50].
  • Quantum Process Tomography (QPT): A method used to fully characterises quantum operations of photonic quantum systems. The idea is to prepare some amount of previously prepared states, perform an operation on these states with quantum operations, and then measure the output states after the operation to obtain reconstruction of the process matrix[14].
@@
  • Noise Spectroscopy: A method that analyzes the overview of the noise frequency spectra in quantum systems by applying certain pulse sequences and measuring the response of a quantum system to those pulse sequences. The resulting noise using these techniques provides information about the dominant sources of noise and their frequency characteristics [12]. “Figure. 7” shows a comparison of noise spectra reconstructed.
Due to the fact that certain traditional approaches will not capture the complete spectrum of complexities associated with noise arising from interactions in photonic quantum systems, there is a need for advanced characterisation techniques designed to enhance understanding of the underlying causes of these types of noise.

### 04 Apr 2026 v3

# Noise Models Impacts and Mitigation Strategies in Photonic Quantum Machine Learning

@@ Abstract
Photonic Quantum Machine Learning (PQML) is an emerging method to implement scalable, energy-efficient quantum information processing by combining photonic quantum computing technologies with machine learning techniques. The features of photonic technologies offer several benefits: room-temperature operation; fast (low delay) processing of signals; and the possibility of representing computations in high-dimensional (Hilbert) spaces. This makes photonic technologies a good candidate for the near-term development of quantum devices. However, noise is still a major limiting factor for the performance, reliability, and scalability of PQML implementations. This review provides a detailed and systematic analysis of the sources of noise that will affect PQML implementations. We will present an overview of the principal photonic quantum computer designs and summarize the many different types of quantum machine learning algorithms that have been successfully implemented using photonic quantum computer architectures such as variational quantum circuits, quantum neural networks, and quantum support vector machines. We identify and categorize the primary sources of noise within photonic quantum systems and how these sources of noise behave algorithm-specifically with respect to degrading the accuracy of learning, unstable training, and slower convergence than expected. Additionally, we review traditional and advanced techniques for characterizing noise and provide an extensive survey of strategies for mitigating the effects of noise on learning performance. Finally, we discuss recent advances that demonstrate PQML’s capability to operate in real-world settings with realistic noise conditions and future obstacles that will challenge the use of PQML as an effective quantum processing platform.
Preprint. This manuscript is currently under review at Discover Quantum Science (Springer Nature).
@@ 1 Introduction
Quantum computing utilizes the principles of quantum mechanics such as superposition and entanglement to solve problems that the classical computer cannot. A single qubit can be represented as a superposition of basis states:
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,$ (1)
where $\alpha,\beta\in\mathbb{C}$ and $|\alpha|^{2}+|\beta|^{2}=1$.
Quantum states can also be represented using density matrices, which are particularly useful for describing noisy systems:
$\rho=|\psi\rangle\langle\psi|$ (2)
For mixed states, the density matrix is given by:
$\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|,$ (3)
where $p_{i}$ are probabilities such that $\sum_{i}p_{i}=1$.
The first-ever ”universal quantum computer” idea was published in 1985 by David Deutsch [15]. In 1994, Peter Shor presents an algorithm that can efficiently find the factors of large numbers, which is later known as Shor’s algorithm [37]. Quantum computation leverages the linear structure of Hilbert space, where quantum states evolve according to unitary transformations, enabling the encoding and manipulation of information in exponentially large state spaces. Entanglement further enables non-classical correlations between quantum systems, which play a crucial role in enhancing computational capabilities beyond classical limits [18]. Quantum computers can revolutionize machine learning by introducing algorithms that run much faster and can process the data more efficiently.
Figure 1: Timeline of major advancements in quantum machine learning.
@@
Quantum Machine Learning (QML) is a discipline that brings together quantum computer science and traditional machine learning. There is no exact date marking the beginning of QML. But Subhash C. Kak published a paper with an idea about a biologically inspired neural network that works on quantum machine learning in 1995 [21]. “Figure. 1” shows the timeline of the major advancements in quantum machine learning. QML seeks to harness the special nature of quantum mechanics to make machine learning more efficient and effective. Quantum computers may have the potential to outperform classical machine learning models for certain tasks by exploiting quantum parallelism and other quantum phenomena. QML currently possesses limitations including noise in quantum systems, limited qubit counts, and the requirement for specialized hardware. However, there is ongoing research into these challenges, and QML has the potential to make great advancements to the machine learning field [16].
@@
Among the various quantum computing platforms, photonic quantum computing has attracted particular interest due to the possibility for operation at room temperature [45] and because photonic systems allow low-latency real-world quantum information processing. In photonic quantum computing, qubits are denoted by photons, which are the basic units of light that can be processed by optical components like beam splitters, phase shifters, mirrors, and others to perform quantum operations. Photonic Quantum Machine Learning (PQML) combines photonic systems with quantum learning algorithms and provides an attractive methodology for potentially scalable, energy-efficient, low-resource, high-fidelity quantum computation [2]. “Figure. 2” illustrates the main components of a photonic quantum computing system.
Figure 2: Schematic overview of a photonic quantum computing system showing optical input processing through a photonic chip and measurement via detectors with classical control electronics.
@@
In this review, we explore the impact of noise in photonic quantum machine learning systems. The upcoming sections of the paper are organized as follows. Section 2 discusses the fundamentals of photonic quantum machine learning. Section 3 identifies and categorizes the various noise sources in photonic quantum systems. Section 4 analyses the impact of noise on quantum machine learning algorithms. Section 5 reviews noise characterization methods. Section 6 explores noise mitigation strategies. Section 7 highlights recent advances and experimental demonstrations. Section 8 provides a discussion, where theoretical frameworks and noise models are considered in the context of recent experimental demonstrations, along with near-term applications and long-term prospects. Finally, Section 9 concludes the review.
@@ 2.1 Photonic Quantum Computing Architectures
There are several architectures for photonic quantum computing, each with distinct noise characteristics and mitigation requirements. But there are only three main architectures that are widely used in photonic quantum computing. These are discrete-variable (DV) quantum computing, continuous-variable (CV) quantum computing, and hybrid quantum computing.
Figure 3: Schematic illustration of a single photon interacting with a beam splitter and being detected at two output ports.
@@
Continuous Variable (CV) Quantum Computing encodes quantum information in continuous degrees of freedom of the infinite-dimensional Hilbert space [27]. Rather than using discrete qubits, CV quantum computers utilize properties of light modes, such as amplitude and phase, to encode and process information. By leveraging Gaussian operations alongside non-Gaussian resources, CV quantum computing allows for universal quantum computation exploiting optical systems[8]. In continuous-variable quantum systems, quantum information is encoded in quadrature operators:
$\hat{x},\hat{p}$ (4)
These operators satisfy the canonical commutation relation:
$\hat{x},\hat{p}]=i\hbar.$ (5)
Noise sources in CV systems include:
@@
Figure 4: Schematic illustration of key components in hybrid quantum systems, including continuous-variable elements such as electromagnetic fields and mechanical oscillators, discrete-variable spin ensembles, phase-space representation via the Wigner function, and homodyne detection.
Hybrid quantum computing amalgamates the two approaches of DV and CV to capitalize on the benefits of both. In hybrid systems we can use discrete qubits encoded in single photons and CV variables for some parts of the measurement or processing. Noise sources in hybrid systems will include DV and CV noise sources, as well as additional complexities produced by the interaction of the discrete and continuous variables [45]. “Figure. 4” illustrates the building blocks of a hybrid quantum system.
@@ 2.2 Quantum Machine Learning Algorithms on Photonic Platforms
Various quantum machine learning algorithms have been adapted to be used on photonic platforms. Some of the prominent algorithms, such as Variational Quantum Circuits (VQC), Quantum Neural Networks (QNN), and Quantum Support Vector Machines (QSVM), have all been successfully implemented in photonic systems [31].So, let’s discuss these algorithms briefly.
Variational Quantum Circuits (VQC) are a powerful framework for quantum machine learning that combines classical optimization techniques and the training of quantum circuits. “Figure. 5” illustrates a basic structure of a VQC. A photonic implementation of VQC can use the huge range of optical components to create tunable quantum circuits, which further allows VQC to successfully train and optimize quantum models [40]. These variational models are particularly sensitive to noise, as the optimization process relies on accurate estimation of expectation values.
A variational quantum circuit is typically represented by a parameterized unitary:
$U(\boldsymbol{\theta})=\prod_{k}e^{-i\theta_{k}H_{k}},$ (6)
where $\boldsymbol{\theta}$ are trainable parameters and $H_{k}$ are Hermitian operators.
The objective of training is to minimize a cost function defined as:
$C(\boldsymbol{\theta})=\langle\psi|U^{\dagger}(\boldsymbol{\theta})OU(\boldsymbol{\theta})|\psi\rangle,$ (7)
where $O$ is an observable.
Quantum Neural Networks (QNN) comprise another category of algorithm that can successfully be implemented on photonic platforms. QNNs use quantum circuits for computations similar to classical neural networks, thus operating in the quantum information domain [24]. A key advantage of photonic QNNs is the inherent parallelism of photonic states, as well as their high-dimensionality, which benefits learning.
A quantum neural network can be expressed as a parameterized quantum transformation:
$y=f\big(U(\boldsymbol{\theta})|x\rangle\big),$ (8)
where $U(\boldsymbol{\theta})$ is a parameterized quantum circuit and $f$ represents measurement and classical post-processing.
Quantum Support Vector Machines (QSVM) are the quantum counterparts of classical support vector machines to be incorporated to handle tasks of classification with the help of quantum states and operations. In the photonic setting, QSVMs can take advantage of the high-dimensional Hilbert space of photons when encoding data and perform kernel evaluations based on quantum interference effects [3].
Quantum support vector machines rely on quantum feature maps and kernel evaluation, defined as:
$K(x_{i},x_{j})=|\langle\phi(x_{i})|\phi(x_{j})\rangle|^{2},$ (9)
where $|\phi(x)\rangle$ represents the quantum feature encoding of input data.
Figure 5: Schematic representation of a variational quantum circuit showing data embedding, parametrized quantum evolution, measurement, and classical post-processing.
@@ 3 Noise Sources in Photonic Quantum Systems
Photonic quantum systems are vulnerable to different sources of noise that degrade any quantum machine learning algorithm that we are conducting [11]. Recognizing and categorizing these sources of noise is paramount to developing suitable noise mitigation strategies. In this section, we discuss the major sources of noise, broken down into two subsections, such as fundamental noise mechanisms and system-specific noise characteristics. These noise processes directly influence the performance of quantum machine learning algorithms, which we analyze in the following section.
The evolution of a quantum state under noise can be described using the Kraus operator formalism:
$\rho^{\prime}=\sum_{k}E_{k}\rho E_{k}^{\dagger},$ (10)
where $E_{k}$ are Kraus operators satisfying $\sum_{k}E_{k}^{\dagger}E_{k}=I$.
Common noise models can be described within this formalism. For example, the depolarizing channel is given by:
$\rho\rightarrow(1-p)\rho+\frac{p}{3}(X\rho X+Y\rho Y+Z\rho Z),$ (11)
where $p$ is the depolarization probability.
Amplitude damping noise can be represented using Kraus operators:
$E_{0}=\begin{bmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{bmatrix},\quad E_{1}=\begin{bmatrix}0&\sqrt{\gamma}\\ 0&0\end{bmatrix}$, (12)
where $\gamma$ is the damping parameter.
These mathematical noise models provide an abstract representation of physical noise processes. In photonic quantum systems, such noise manifests through practical mechanisms such as photon loss, phase noise, and detector inefficiencies.
@@ 3.1 Fundamental Noise Mechanisms
While photons are excellent carriers of quantum information, they are still subject to various kinds of noise during generation, transmission, and detection, which can degrade the quality of quantum states and diminish the reliability of practitioners’ quantum operations [19]. Below, we identify the main types of noise in photonic quantum systems and how these types of noise affect the overall performance of a system.
@@
Photon loss can have a detrimental effect on quantum machine learning algorithms. As an example, in variational quantum circuits, photon loss can result in an erroneous adjustment of parameters while training and exhibit poor model performance [11]. With respect to quantum neural networks, photon loss can sever the flow of information in the network, degrading learning capabilities [1].
@@ 3.2 System-Specific Noise Characteristics
The other type of noise source is related to the specific design of a particular photonic quantum device, and the way that the device itself is manufactured, like architectures and components, will also affect how the specific types of noise behave. In this subsection, many examples of photonic quantum device-specific noise have been provided.
  • Mode Mismatch: Mismatched optical modes often cause reduced visibility of interference, which lowers the values of interference[42]. Also increased error rates on quantum algorithms that rely on interference like boson sampling, quantum walks, etc.
@@
As quantum machine learning models rely heavily on the presence of high-visibility interference, mode mismatch is one of the most detrimental sources of noise in photonic implementations. The next section will provide a thorough discussion of how mode mismatch and other types of noise represent one of the major sources of impact on photonic implementations from a practical standpoint.
@@ 4 Impact of Noise on Quantum Machine Learning Algorithms
The expectation value of an observable is given by:
$\langle O\rangle=\mathrm{Tr}(\rho O),$ (13)
where $\mathrm{Tr}$ denotes the trace operation. Noise modifies the quantum state $\rho$, thereby affecting the measured expectation values and degrading learning performance.
@@ 4.1 Algorithm-Specific Noise Sensitivity
Noise can have a significant impact on the operation of VQCs and QNNs. Due to their reliance on the accurate measurement of expectation values and calculation of gradients, variational quantum circuits (VQCs) are especially vulnerable to the presence of noise [14]. The impact of noise on VQC can be affected in different ways, such as:
  • Gradient Vanishing/Exploding: Noise can lead to ”vanishing” or ”exploding” gradients during the optimization phase, making it hard to develop a good quantum model because the gradient is either close to zero or is extremely large. [30]
@@
QNNs operate using quantum mechanics and can be easily impacted by noise during their operation. Therefore, noise may impact the ability of QNNs to learn and adapt. Similarly, feedforward structures, convolutional, and recurrent QNNs will each experience issues with noise at various levels. So we have to use different strategies to mitigate these noise effects.
@@ 4.2 Noise-Resilient Algorithm Design
Noise-resilient algorithm design focuses on creating quantum machine learning algorithms that can withstand the presence of noise in photonic quantum systems. According to recent research, we can get mainly two types: hardware-aware compilation and algorithmic error suppression. “Figure. LABEL:fig6” illustrates the concept of quantum error mitigation (QEM), which aims to design noise-resilient algorithms.
Hardware-aware compiling tools for modern quantum circuits take advantage of the specific types of noise present in the underlying hardware to create quantum circuits that are as close to optimal as possible. Specifically, the optimized circuit designs reduce the number of overall errors based on the type of noise present in the photonic platform used. The optimizations include reducing circuit depths and limiting the number of entangling operations within a particular circuit design to reduce the negative effect of photon loss and decoherence.
@@ 5.1 Traditional Noise Characterization Approaches
  • Quantum Process Tomography (QPT): A method used to fully characterize quantum operations of photonic quantum systems. The idea is to prepare some amount of previously prepared states, perform an operation on these states with quantum operations, and then measure the output states after the operation to obtain a reconstruction of the process matrix[12].
@@
  • Noise Spectroscopy: A method that analyzes the overview of the noise frequency spectra in quantum systems by applying certain pulse sequences and measuring the response of a quantum system to those pulse sequences. The resulting noise using these techniques provides information about the dominant sources of noise and their frequency characteristics [10]. “Figure. LABEL:fig7” shows a comparison of noise spectra reconstructed.
Due to the fact that certain traditional approaches will not capture the complete spectrum of complexities associated with noise arising from interactions in photonic quantum systems, there is a need for advanced characterization techniques designed to enhance understanding of the underlying causes of these types of noise.